# Cafe-Mpc: A Cascaded-Fidelity Model Predictive Control Framework with Tuning-Free Whole-Body Control

### 09 Mar 2024 v2

@@ Abstract
This work introduces an optimization-based locomotion control framework for on-the-fly synthesis of complex dynamic maneuvers. At the core of the proposed framework is a cascaded-fidelity model predictive controller (Cafe-Mpc). Cafe-Mpc strategically relaxes the planning problem along the prediction horizon (i.e., with descending model fidelity, increasingly coarse time steps, and relaxed constraints) for computational and performance gains. This problem is numerically solved with an efficient customized multiple-shooting iLQR (MS-iLQR) solver that is tailored for hybrid systems. The action-value function from Cafe-Mpc is then used as the basis for a new value-function-based whole-body control (VWBC) technique that avoids additional tuning for the WBC. In this respect, the proposed framework unifies whole-body MPC and more conventional whole-body quadratic programming (QP), which have been treated as separate components in previous works. We study the effects of the cascaded relaxations in Cafe-Mpc on the tracking performance and required computation time. We also show that the Cafe-Mpc , if configured appropriately, advances the performance of whole-body MPC without necessarily increasing computational cost. Further, we show the superior performance of the proposed VWBC over the Ricatti feedback controller in terms of constraint handling. The proposed framework enables accomplishing for the first time gymnastic-style running barrel roll on the MIT Mini Cheetah. Video link.
@@ I Introduction
The proposed framework is built upon model predictive control (MPC), a method of producing control inputs by predicting the future behavior of the robot as part of solving a trajectory optimization (TO) problem. MPC accounts for model uncertainties and external disturbances by frequently adapting the predicted plan to the most recent state. Whole-body MPC produces a high-fidelity plan that is coherent with the full robot dynamics. However, it is known to have large computational burdens, prohibiting the use of a long prediction horizon or fast update frequency. For this reason, the capability of this approach to control highly dynamic motions on robot hardware has not yet been demonstrated [15, 16, 17]. A local feedback Ricatti controller [15] that runs at much higher frequency can be used to account for the policy delay and the slow MPC update. This feedback controller, however, is not guaranteed to respect the constraints in the whole-body MPC. A notable strategy to alleviate the computational burden is to instead use a simplified template model, thus template MPC. A low-level controller that converts the low-dimensional plan to whole-body commands is needed, often an inverse-dynamics quadratic program (QP). This strategy is thus far the most popular MPC approach, and has demonstrated great success for legged locomotion [18, 11, 5, 19, 20, 21, 22]. However, the omitted model details and the possibly added artificial constraints may result in infeasible or over-conservative target plans [23]. As a result, nontrivial tuning of the whole-body QP is often needed. The proposed framework introduces a novel approach that accelerates the whole-body MPC, and a novel low-level controller that integrates the Ricatti controller and the whole-body QP.
@@ II-C Whole-Body Control
The use of a Ricatti controller is a less-known but powerful WBC technique. The feedback gains can be computed either via DDP/iLQR [66, 79, 15], or via multiple-shooting based numerical optimization [5], though the ideas are similar where the Difference Ricatti equations are computed. The Ricatti controller can either be directly used if the MPC reasons about whole-body dynamics [15], or an additional layer is needed to produce the whole-body actuation if the MPC reasons about simplified models [66, 79]. A known problem of this approach is that the feedback control does not necessarily satisfy constraints, which can interfere with robot stability, for instance, when the friction cone constraints are violated.
The value-based WBC (VWBC) proposed in this work overcomes the issues of the conventional OSC-QP and the Ricatti controller. The VWBC is not significantly different from the OSC-QP in terms of formulation. The novelty is that it employs an objective function that considers the long-term cost-to-go (value function). This value function is a proxy of the long-term stability, and is readily available from the top-level Cafe-Mpc, thus leaving the VWBC free of additional tuning. The VWBC generalizes the Ricatti controller to include constraints, thereby unifying whole-body MPC with conventional WBC.
@@ III-A MS-DDP for Single-Phase TO
Giftthaler[75] proposes a flexible formulation that allows a subset of $\mathbf{X}$ to be independent while keeping the rest dependent variables. These independent variables are called shooting states, the indices of which are collected to a set $\mathbb{{M}}$. These dependent variables are called roll-out states, the indices of which are collected to the complementary set $\tilde{\mathbb{{M}}}$. If $k\in\mathbb{{M}}$, then $\bar{\mathbf{x}}_{k}$ is seeded from an initial guess. Otherwise, it is overwritten by the simulated state $\mathbf{f}(\bar{\mathbf{x}}_{k},\bar{\mathbf{u}}_{k})$. MS-DDP iteratively improves $(\bar{\mathbf{X}},\bar{\mathbf{U}})$ until the defects are sufficiently small and the cost function is minimized. At each iteration, it performs a backward sweep followed by a forward sweep within a line search process.
@@ IV-A1 Cost Function
where the operation $\delta\cdot$ denotes the difference between a variable and its reference, $\mathbf{S}$ is a diagonal matrix of $\{0,1\}$ elements indicating swing leg status. The terminal cost is similarly defined but without the last term.
@@ IV-A2 Dynamics
$\begin{bmatrix}\mathbf{M}&\mathbf{J}_{c}^{T}\\ \mathbf{J}_{c}&\mathbf{0}\end{bmatrix}\begin{bmatrix}\ddot{\mathbf{q}}\\ -\mathbf{F}_{c}\end{bmatrix}=\begin{bmatrix}\mathbf{S}^{T}\boldsymbol{\tau}-\mathbf{n}(\mathbf{q},\dot{\mathbf{q}})\\ -\dot{\mathbf{J}}_{c}\dot{\mathbf{q}}+\alpha\mathbf{J}_{c}\dot{\mathbf{q}}\end{bmatrix}.$ (21)
@@
$\mathbf{x}^{[i+1]}_{0}=\begin{bmatrix}{(\mathbf{q}^{[i]^{-}})}^{\top},{(\dot{\mathbf{q}}^{[i]^{+}})}^{\top}\end{bmatrix}^{\top}:=\mathbf{P}_{\text{w}}^{[i]}(\mathbf{x}^{[i]}_{0}).$ (24)
@@ IV-B1 Cost Function
where $\mathbf{W}_{\theta,c}$ are the components in $\mathbf{W}_{q}$ corresponding to $\theta$ and $\mathbf{c}$, and $\mathbf{W}_{\dot{\theta},\dot{c}}$ are the components in $\mathbf{W}_{\dot{q}}$ corresponding to $\dot{\boldsymbol{\theta}}$ and $\dot{\mathbf{c}}$, and $\tilde{\mathbf{S}}=\mathbf{1}_{4}-\mathbf{S}$ is the contact status matrix where $\mathbf{1}_{4}$ is the 4$\times$4 identity matrix and $\mathbf{S}$ denotes swing status as defined in Eq. (18). The terminal cost is similarly defined but excludes the regularization term.
@@ V Value-Based Whole-Body Controller
where $\mathbf{K}=-{\mathbf{Q}}^{-1}_{\mathbf{u}\mathbf{u}}\mathbf{Q}_{\mathbf{u}\mathbf{x}}$ as in Eq. (7). This approach is known as the Ricatti feedback control. It smooths the actuation torque in between MPC time steps, enables improved stability, and allows for slower MPC update [66, 15]. Despite these benefits, a well-known problem with (33) is that the closed-loop trajectory does not necessarily satisfy certain physical constraints. This problem can be critical to the robot stability, for instance, the robot can fall when friction cone constraints are violated.
@@ V-A Value-Function Based WBC
The value-function-based WBC (VWBC) proposed in this work embeds the Ricatti feedback controller (33) within a conventional WBC-QP formulation, thereby unifying the WBC and the Ricatti feedback control. The VWBC produces a whole-body control command that is close to the Ricatti controller, while at the same time satisfying all necessary physical constraints. The VWBC is formulated as a QP
@@
The VWBC (V-A) is a generalization of the Ricatti controller (7). This can be seen by removing all the constraints in (V-A). The problem (V-A) then becomes an unconstrained optimization problem, the solution of which is Eq. (7). Thus, the VWBC seeks a control signal that is the closest to the Ricatti controller (7) but with all necessary constraints satisfied. In other words, the VWBC can be considered as the Ricatti controller disguised in a QP, thus enjoying the benefit of feedback stabilization, while preventing the resulting solution from being too aggressive to violate the constraints. The VWBC is similar in spirit as [77, 80], except that the value functions in [77, 80] are obtained via LQR designed for simplified models.
@@
The QP (V-A) can further be warm started using solutions of Cafe-Mpc, and thus requiring fewer solver iterations. The whole-body plan of Cafe-Mpc incorporates $\boldsymbol{\tau}_{k}^{*},\mathbf{q}_{k}^{*},\dot{\mathbf{q}}_{k}^{*},\mathbf{F}_{c,k}^{*}$. We obtain $\ddot{\mathbf{q}}^{*}=(\dot{\mathbf{q}}_{k+1}^{*}-\dot{\mathbf{q}}_{k}^{*})/d\text{t}_{\text{w}}$, and use $\boldsymbol{\tau}_{k}^{*},\ddot{\mathbf{q}}^{*},\mathbf{F}_{c,k}^{*}$ as an initial guess for the QP (V-A). As will be shown in Section 9, only one iteration is needed most of the time for a bounding gait with an active set method.
@@ VI-C Engineering Details
While the VWBC runs on the robot hardware at 500 Hz, the Cafe-Mpc is executed on a separate computer at 33 to 50 Hz depending on the motion task. The communication between them is carried out via LCM[104]. To account for the policy lag due to MPC solve time and communication latency, a sub-trajectory of size six in the front of the MPC solution is sent to the low-level VWBC controller. The VWBC then finds a solution that is closest to the current time. In addition, since the VWBC plays a role of Ricatti feedback controller and executes at a significantly higher rate than Cafe-Mpc, it further also minimizes the effect of policy lag.
@@ VII-A2 Hardware Results
Figure 10: Time-series snapshots of Mini Cheetah bounding under forward velocity strong disturbances (1.5 m/s) in simulation using Cafe-Mpc, and different whole-body control schemes. Top: Ricatti feedback controller. Bottom: proposed VWBC.
@@ VII-B VWBC
In this section, we study the performance of the VWBC, and compare it with the performance of the Ricatti controller. To remind the readers of the differences, the Ricatti controller controls robots with the feedforward-feedback controller (7) resulting from the MS-iLQR solver. The VWBC solves a value-based QP (V-A) that embeds (7), but with all necessary constraints satisfied. Toward that regard, we benchmark the differences with a disturbance rejection test, and are mainly concerned with constraint satisfaction and disturbance recovery. We use qpOASES [105] as the QP solver. The whole-body plan of Cafe-Mpc offers an initial guess for the valued-based QP without additional computational cost. We compare the QP solve times with and without using this initial guess.
@@ VII-B1 Disturbance Recovery
The disturbance rejection test was conducted with the same bounding gait as in Section IV. We change the body velocity by a total of 1.5 m/s over 50 ms, which corresponds to an effective external for of 270 N over the period. Figure. 9 depicts the joint torques of the front right and back right legs as well as their upper and lower limits. It demonstrates that the Ricatti controller sometimes violates the torque constraints, especially after the push disturbance, whereas the VWBC consistently satisfies the torque constraints. One may argue that a simple clamping technique can work here, and question the need for the proposed technique. This argument is true in this case, and clamping torque is sufficient to prevent control saturation. However, the VWBC provides a systematic way of balancing optimality (local Q function) and constraint satisfaction. Further, some other constraints such as friction that cannot be clamped are important to robot stability. To demonstrate, Fig. 10 shows time-series snapshots of the robot after the push disturbance. With the Ricatti controller, the robot has a non-trivial slipping that eventually disablizes the robot. With the VWBC, the slipping is slight, and the robot quickly recovers a stable contact in one gait cycle.
@@ VIII-A Discussions and Future Works
The VWBC disguises the Ricatti feedback controller in a QP problem. Results in Section VII-B demonstrate that the VWBC enjoys the benefits of Ricatti feedback closing the control loop at a higher rate, while at the same time satisfying necessary constraints. This feature allows the high-level Cafe-Mpc to be solved at coarse precision. One limitation of VWBC is that the foot contacts are currently specified using a known schedule, which can cause instability problems in the presence of a large contact mismatch. Future work would seed VWBC with contact detection [107].

### 14 Mar 2024 v3

@@ Abstract
This work introduces an optimization-based locomotion control framework for on-the-fly synthesis of complex dynamic maneuvers. At the core of the proposed framework is a cascaded-fidelity model predictive controller (Cafe-Mpc). Cafe-Mpc strategically relaxes the planning problem along the prediction horizon (i.e., with descending model fidelity, increasingly coarse time steps, and relaxed constraints) for computational and performance gains. This problem is numerically solved with an efficient customized multiple-shooting iLQR (MS-iLQR) solver that is tailored for hybrid systems. The action-value function from Cafe-Mpc is then used as the basis for a new value-function-based whole-body control (VWBC) technique that avoids additional tuning for the WBC. In this respect, the proposed framework unifies whole-body MPC and more conventional whole-body quadratic programming (QP), which have been treated as separate components in previous works. We study the effects of the cascaded relaxations in Cafe-Mpc on the tracking performance and required computation time. We also show that the Cafe-Mpc , if configured appropriately, advances the performance of whole-body MPC without necessarily increasing computational cost. Further, we show the superior performance of the proposed VWBC over the Riccati feedback controller in terms of constraint handling. The proposed framework enables accomplishing for the first time gymnastic-style running barrel roll on the MIT Mini Cheetah. Video link.
@@ I Introduction
The proposed framework is built upon model predictive control (MPC), a method of producing control inputs by predicting the future behavior of the robot as part of solving a trajectory optimization (TO) problem. MPC accounts for model uncertainties and external disturbances by frequently adapting the predicted plan to the most recent state. Whole-body MPC produces a high-fidelity plan that is coherent with the full robot dynamics. However, it is known to have large computational burdens, prohibiting the use of a long prediction horizon or fast update frequency. For this reason, the capability of this approach to control highly dynamic motions on robot hardware has not yet been demonstrated [15, 16, 17]. A local feedback Riccati controller [15] that runs at much higher frequency can be used to account for the policy delay and the slow MPC update. This feedback controller, however, is not guaranteed to respect the constraints in the whole-body MPC. A notable strategy to alleviate the computational burden is to instead use a simplified template model, thus template MPC. A low-level controller that converts the low-dimensional plan to whole-body commands is needed, often an inverse-dynamics quadratic program (QP). This strategy is thus far the most popular MPC approach, and has demonstrated great success for legged locomotion [18, 11, 5, 19, 20, 21, 22]. However, the omitted model details and the possibly added artificial constraints may result in infeasible or over-conservative target plans [23]. As a result, nontrivial tuning of the whole-body QP is often needed. The proposed framework introduces a novel approach that accelerates the whole-body MPC, and a novel low-level controller that integrates the Riccati controller and the whole-body QP.
@@ II-C Whole-Body Control
The use of a Riccati controller is a less-known but powerful WBC technique. The feedback gains can be computed either via DDP/iLQR [66, 79, 15], or via multiple-shooting based numerical optimization [5], though the ideas are similar where the Difference Riccati equations are computed. The Riccati controller can either be directly used if the MPC reasons about whole-body dynamics [15], or an additional layer is needed to produce the whole-body actuation if the MPC reasons about simplified models [66, 79]. A known problem of this approach is that the feedback control does not necessarily satisfy constraints, which can interfere with robot stability, for instance, when the friction cone constraints are violated.
The value-based WBC (VWBC) proposed in this work overcomes the issues of the conventional OSC-QP and the Riccati controller. The VWBC is not significantly different from the OSC-QP in terms of formulation. The novelty is that it employs an objective function that considers the long-term cost-to-go (value function). This value function is a proxy of the long-term stability, and is readily available from the top-level Cafe-Mpc, thus leaving the VWBC free of additional tuning. The VWBC generalizes the Riccati controller to include constraints, thereby unifying whole-body MPC with conventional WBC.
@@ III-A MS-DDP for Single-Phase TO
Giftthaler[75] proposes a flexible formulation that allows a subset of $\mathbf{X}$ to be independent while keeping the rest dependent variables. These independent variables are called shooting states, the indices of which are collected to a set $\mathbb{{M}}$. These dependent variables are called roll-out states, the indices of which are collected to the complementary set $\tilde{\mathbb{{M}}}$. If $k\in\mathbb{{M}}$, then $\bar{\mathbf{x}}_{k}$ is seeded from an initial guess. Otherwise, it is overwritten by the simulated state $\mathbf{f}(\bar{\mathbf{x}}_{k-1},\bar{\mathbf{u}}_{k-1})$. MS-DDP iteratively improves $(\bar{\mathbf{X}},\bar{\mathbf{U}})$ until the defects are sufficiently small and the cost function is minimized. At each iteration, it performs a backward sweep followed by a forward sweep within a line search process.
@@ IV-A1 Cost Function
where the operation $\delta\cdot$ denotes the difference between a variable and its reference. Let $s_{j}$ denote the swing status of the $j^{\text{th}}$ leg, then $\mathbf{S}\in\mathbb{{R}}^{12\times 12}$ is a diagonal matrix whose diagonal concatenates vectors $[s_{j},s_{j},s_{j}]$ of all four legs. The terminal cost is similarly defined but without the last term.
@@ IV-A2 Dynamics
$\begin{bmatrix}\mathbf{M}&\mathbf{J}_{c}^{T}\\ \mathbf{J}_{c}&\mathbf{0}\end{bmatrix}\begin{bmatrix}\ddot{\mathbf{q}}\\ -\mathbf{F}_{c}\end{bmatrix}=\begin{bmatrix}\mathbf{B}_{\tau}^{T}\boldsymbol{\tau}-\mathbf{n}(\mathbf{q},\dot{\mathbf{q}})\\ -\dot{\mathbf{J}}_{c}\dot{\mathbf{q}}+\alpha\mathbf{J}_{c}\dot{\mathbf{q}}\end{bmatrix}.$ (21)
@@
$\mathbf{x}^{[i+1]}_{0}=\begin{bmatrix}{(\mathbf{q}^{[i]^{-}})}^{\top},{(\dot{\mathbf{q}}^{[i]^{+}})}^{\top}\end{bmatrix}^{\top}:=\mathbf{P}_{\text{w}}^{[i]}(\mathbf{x}^{[i]}_{N}).$ (24)
@@ IV-B1 Cost Function
where $\mathbf{W}_{\theta,c}$ are the components in $\mathbf{W}_{q}$ corresponding to $\theta$ and $\mathbf{c}$, and $\mathbf{W}_{\dot{\theta},\dot{c}}$ are the components in $\mathbf{W}_{\dot{q}}$ corresponding to $\dot{\boldsymbol{\theta}}$ and $\dot{\mathbf{c}}$, and $\tilde{\mathbf{S}}=\mathbf{1}_{12}-\mathbf{S}$ is the contact status matrix where $\mathbf{1}_{12}\in\mathbb{{R}}^{12\times 12}$ is an identity matrix and $\mathbf{S}$ denotes swing status as defined in Eq. (18). The terminal cost is similarly defined but excludes the regularization term.
@@ V Value-Based Whole-Body Controller
where $\mathbf{K}=-{\mathbf{Q}}^{-1}_{\mathbf{u}\mathbf{u}}\mathbf{Q}_{\mathbf{u}\mathbf{x}}$ as in Eq. (7). This approach is known as the Riccati feedback control. It smooths the actuation torque in between MPC time steps, enables improved stability, and allows for slower MPC update [66, 15]. Despite these benefits, a well-known problem with (33) is that the closed-loop trajectory does not necessarily satisfy certain physical constraints. This problem can be critical to the robot stability, for instance, the robot can fall when friction cone constraints are violated.
@@ V-A Value-Function Based WBC
The value-function-based WBC (VWBC) proposed in this work embeds the Riccati feedback controller (33) within a conventional WBC-QP formulation, thereby unifying the WBC and the Riccati feedback control. The VWBC produces a whole-body control command that is close to the Riccati controller, while at the same time satisfying all necessary physical constraints. The VWBC is formulated as a QP
@@
The VWBC (V-A) is a generalization of the Riccati controller (7). This can be seen by removing all the constraints in (V-A). The problem (V-A) then becomes an unconstrained optimization problem, the solution of which is Eq. (7). Thus, the VWBC seeks a control signal that is the closest to the Riccati controller (7) but with all necessary constraints satisfied. In other words, the VWBC can be considered as the Riccati controller disguised in a QP, thus enjoying the benefit of feedback stabilization, while preventing the resulting solution from being too aggressive to violate the constraints. The VWBC is similar in spirit as [77, 80], except that the value functions in [77, 80] are obtained via LQR designed for simplified models.
@@
The QP (V-A) can further be warm started using solutions of Cafe-Mpc, and thus requiring fewer solver iterations. The whole-body plan of Cafe-Mpc incorporates $\boldsymbol{\tau}_{k}^{*},\mathbf{q}_{k}^{*},\dot{\mathbf{q}}_{k}^{*},\mathbf{F}_{c,k}^{*}$. We obtain $\ddot{\mathbf{q}}_{k}^{*}=(\dot{\mathbf{q}}_{k+1}^{*}-\dot{\mathbf{q}}_{k}^{*})/d\text{t}_{\text{w}}$, and use $\boldsymbol{\tau}_{k}^{*},\ddot{\mathbf{q}}_{k}^{*},\mathbf{F}_{c,k}^{*}$ as an initial guess for the QP (V-A). As will be shown in Section 9, only one iteration is needed most of the time for a bounding gait with an active set method.
@@ VI-C Engineering Details
While the VWBC runs on the robot hardware at 500 Hz, the Cafe-Mpc is executed on a separate computer at 33 to 50 Hz depending on the motion task. The communication between them is carried out via LCM[104]. To account for the policy lag due to MPC solve time and communication latency, a sub-trajectory of size six in the front of the MPC solution is sent to the low-level VWBC controller. The VWBC then finds a solution that is closest to the current time. In addition, since the VWBC plays a role of Riccati feedback controller and executes at a significantly higher rate than Cafe-Mpc, it further also minimizes the effect of policy lag.
@@ VII-A2 Hardware Results
Figure 10: Time-series snapshots of Mini Cheetah bounding under forward velocity strong disturbances (1.5 m/s) in simulation using Cafe-Mpc, and different whole-body control schemes. Top: Riccati feedback controller. Bottom: proposed VWBC.
@@ VII-B VWBC
In this section, we study the performance of the VWBC, and compare it with the performance of the Riccati controller. To remind the readers of the differences, the Riccati controller controls robots with the feedforward-feedback controller (7) resulting from the MS-iLQR solver. The VWBC solves a value-based QP (V-A) that embeds (7), but with all necessary constraints satisfied. Toward that regard, we benchmark the differences with a disturbance rejection test, and are mainly concerned with constraint satisfaction and disturbance recovery. We use qpOASES [105] as the QP solver. The whole-body plan of Cafe-Mpc offers an initial guess for the valued-based QP without additional computational cost. We compare the QP solve times with and without using this initial guess.
@@ VII-B1 Disturbance Recovery
The disturbance rejection test was conducted with the same bounding gait as in Section IV. We change the body velocity by a total of 1.5 m/s over 50 ms, which corresponds to an effective external for of 270 N over the period. Figure. 9 depicts the joint torques of the front right and back right legs as well as their upper and lower limits. It demonstrates that the Riccati controller sometimes violates the torque constraints, especially after the push disturbance, whereas the VWBC consistently satisfies the torque constraints. One may argue that a simple clamping technique can work here, and question the need for the proposed technique. This argument is true in this case, and clamping torque is sufficient to prevent control saturation. However, the VWBC provides a systematic way of balancing optimality (local Q function) and constraint satisfaction. Further, some other constraints such as friction that cannot be clamped are important to robot stability. To demonstrate, Fig. 10 shows time-series snapshots of the robot after the push disturbance. With the Riccati controller, the robot has a non-trivial slipping that eventually disablizes the robot. With the VWBC, the slipping is slight, and the robot quickly recovers a stable contact in one gait cycle.
@@ VIII-A Discussions and Future Works
The VWBC disguises the Riccati feedback controller in a QP problem. Results in Section VII-B demonstrate that the VWBC enjoys the benefits of Riccati feedback closing the control loop at a higher rate, while at the same time satisfying necessary constraints. This feature allows the high-level Cafe-Mpc to be solved at coarse precision. One limitation of VWBC is that the foot contacts are currently specified using a known schedule, which can cause instability problems in the presence of a large contact mismatch. Future work would seed VWBC with contact detection [107].